# UVIP: Model-Free Approach to Evaluate Reinforcement Learning Algorithms

### v4

@@ Abstract
Policy evaluation is an important instrument for the comparison of different algorithms in Reinforcement Learning (RL). Yet even a precise knowledge of the value function $V^{\pi}$ corresponding to a policy $\pi$ does not provide reliable information on how far is the policy $\pi$ from the optimal one. We present a novel model-free upper value iteration procedure (UVIP) that allows us to estimate the suboptimality gap $V^{\star}(x)-V^{\pi}(x)$ from above and to construct confidence intervals for $V^{\star}$. Our approach relies on upper bounds to the solution of the Bellman optimality equation via martingale approach. We provide theoretical guarantees for UVIP under general assumptions and illustrate its performance on a number of benchmark RL problems.
@@ 1 Introduction
The key objective of Reinforcement Learning (RL) is to learn an optimal agent’s behaviour in an unknown environment. A natural performance metric is given by the value function $V^{\pi}$ which is expected total reward of the agent following $\pi$. There are efficient algorithms to evaluate this quantity, e.g. temporal difference methods Sutton (1988), Tsitsiklis & Van Roy (1997). Unfortunately, even a precise knowledge of $V^{\pi}$ does not provide reliable information on how far is the policy $\pi$ from the optimal one. To address this issue a popular quality measure is the regret of the algorithm that is the difference between the total sum of rewards accumulated when following the optimal policy and the sum of rewards obtained when following the current policy $\pi$ (see e.g. Jaksch et al. (2010)). In the setting of finite state- and action space Markov Decision Processes (MDP) there is a variety of regret bounds for popular RL algorithms like Q-learning Jin et al. (2018), optimistic value iteration Azar et al. (2017), and many others.
Unfortunately, regret bounds beyond the discrete setup are much less common in the literature. Even more crucial drawback of the regret-based comparison is that regret bounds are typically pessimistic and rely on the unknown quantities of the underlying MDP’s. A simpler, but related, quantity is the suboptimality gap (policy error) $\Delta_{\pi}(x)\doteq V^{\star}(x)-V^{\pi}(x)$. Since we do not know $V^{\star}$, the suboptimality gap can not be calculated directly. There is a vast amount of literature devoted to theoretical guarantees for $\Delta_{\pi}(x)$, see e.g. Antos et al. (2007), Szepesvári (2010), Pires & Szepesvári (2016) and references therein. However, these bounds share the same drawbacks as the regret bounds. Moreover, known bounds does not apply to the general policy $\pi$ and depends heavily on the particular algorithm which produced it. For instance, in Approximate Policy Iteration (API, Bertsekas & Tsitsiklis (1996)) all existing bounds for $\Delta_{\pi}(x)$ depend on the one-step error induced by the approximation of the action-value function. This one-step error is difficult to quantify since it depends on the unknown smoothness properties of the action-value function. Similarly, in policy gradient methods (see e.g. Sutton & Barto (2018)), there is always an approximation error due to the choice of family of policies that can be hardly quantified. The approach based on the policy optimism principle (see Efroni et al. (2019)) suggests to initialise the value iteration algorithm using an upper bound (optimistic value) for $V^{\star}$, yielding a sequence of upper bounds converging to $V^{\star}$. However this approach is tailored to finite state- and action space MDPs and is not applicable to evaluate the quality of the general policy $\pi$. To summarise, such bounds can not be used to construct tight model-free confidence bounds for $V^{\star}.$
In this paper we are interested in deriving agnostic (model independent) bounds for the policy error using the concept of upper solutions to the Bellman optimality equation. Our approach is substantially different from the ones known in literature as it can be used to evaluate performance bounds for arbitrary given policy $\pi$. The concept of upper solutions is closely related to martingale duality in optimal control and information relaxation approach, see Belomestny & Schoenmakers (2018), Rogers (2007) and references therein. This idea has been successfully used in the recent paper Shar & Jiang (2020). This work proposes to use the duality approach to improve the performance of Q-learning algorithm in finite horizon MDP through the use of “lookahead” upper and lower bounds. Compared to Shar & Jiang (2020), our approach is not restricted to the Q-learning. The concept of upper solutions has also a connection to distributional RL, as it can be formulated pathwise or using distributional Bellman operator, see e.g. Lyle et al. (2019). A further study of this connection is a promising future research area.
@@
  • We propose a novel approach to construct model free confidence bounds for the optimal value function $V^{\star}$ based on a notion of upper solutions.
@@
  • We study convergence properties of the approximate UVIP in the case of general state and action spaces. In particular, we show that the variance of the resulting upper bound is small if $\pi$ is close to $\pi^{*}$ leading to the tight confidence bounds for $V^{\star}$.
The paper is organized as follows. First, in section 2, we briefly recall main concepts related with the MDPs, and introduce some notations. Then in sections 3 and 4 we introduce the framework of UVIP and discuss its basic properties. In section 5 we perform theoretical study of the approximate UVIP. Numerical results are collected in section 6. Section 7 concludes the paper. Section A in appendix is devoted to the proof of main theoretical results.
Notations and definitions
For $N\in\mathbb{N}$ we define $[N]\doteq\{1,\ldots,N\}$. Let us denote the space of bounded measurable functions with domain $\mathsf{X}$ by $\mathcal{B}(\mathsf{X})$ equipped with the norm $\|f\|_{\mathsf{X}}=\sup_{x\in\mathsf{X}}|f(x)|$ for any $f\in\mathcal{B}(\mathsf{X}).$ In what follows, whenever a norm is uniquely identifiable from its argument, we will drop the index of the norm denoting the underlying space. We denote by ${\rm P}^{a}$ an $\mathcal{B}(\mathsf{X})\to\mathcal{B}(\mathsf{X})$ operator defined by $({\rm P}^{a}V)(x)=\int V(x^{\prime}){\rm P}^{a}(dx^{\prime}|x)$. For an arbitrary metric space $(\mathcal{X},\rho_{\mathcal{X}})$ and function $f:\mathcal{X}\to\mathbb{R}$ we define by $\mathrm{Lip}_{\rho_{\mathcal{X}}}(f)\doteq\sup_{x\neq y}|f(x)-f(y)|/\rho_{\mathcal{X}}(x,y)$.
@@ 2 Preliminary
A Markov Decision Process (MDP) is a tuple $(\mathsf{X},\mathsf{A},\mathcal{P},r)$, where $\mathsf{X}$ is the state space, $\mathsf{A}$ is the action space, $\mathcal{P}=({\rm P}^{a})_{a\in\mathsf{A}}$ is the transition probability kernel and $r=(r^{a})_{a\in\mathsf{A}}$ is the reward function. For each state $x\in\mathsf{X}$ and action $a\in\mathsf{A}$, ${\rm P}^{a}(\cdot|x)$ stands for a distribution over the states in $\mathsf{X}$, that is, the distribution over the next states given that action $a$ is taken in the state $x$. For each action $a\in\mathsf{A}$ and state $x\in\mathsf{X}$, $r^{a}(x)$ gives a reward received when action $a$ is taken in state $x$. An MDP describes the interaction of an agent and its environment. When an action $A_{t}\in\mathsf{A}$ at time $t$ is chosen by the agent, the state $X_{t}$ transitioned to $X_{t+1}\sim{\rm P}^{A_{t}}(\cdot|X_{t})$. The agent’s goal is to maximize the expected total discounted reward, $\mathsf{E}[\sum_{t=0}^{\infty}\gamma^{t}r^{A_{t}}(X_{t})]$, where $10<\gamma<$ 1 is the discount factor. A rule describing the way an agent acts given its past actions and observations is called a policy. The value function of a policy $\pi$ in a state $x\in\mathsf{X}$, denoted by $V^{\pi}(x)$, is $V^{\pi}(x)=\mathsf{E}[\sum_{t=0}^{\infty}\gamma^{t}r^{A_{t}}(X_{t})|X_{0}=x]$, that is, the expected total discounted reward when the initial state ($X_{0}=x$) assuming the agent follows the policy $\pi$. Similarly, we define the action-value function $Q^{\pi}(x,a)=\mathsf{E}[\sum_{t=0}^{\infty}\gamma^{t}r^{A_{t}}(X_{t})|X_{0}=x,A_{0}=a]$. An optimal policy is one that achieves the maximum possible value amongst all policies in each state $x\in\mathsf{X}$. The optimal value for state $x$ is denoted by $V^{\star}(x)$. A deterministic Markov policy can be identified with a map $\pi:\mathsf{X}\to\mathsf{A}$, and the space of measurable deterministic Markov policies will be denoted by $\Pi$. When, in addition, the reward function is bounded, which we assume from now on, all the value functions are bounded and one can always find a deterministic Markov policy that is optimal Puterman (2014). We also define a greedy policy w.r.t. action-value function $Q(x,a)$, which is a deterministic policy $\pi(x)\in\operatorname*{argmax}_{a\in\mathsf{A}}Q(x,a)$. The Bellman return operator w.r.t. ${\rm P}$, $\mathsf{T}_{{\rm P}}:\mathcal{B}(\mathsf{X})\to\mathcal{B}(\mathsf{X}\times\mathsf{A})$, is defined by $(\mathsf{T}_{{\rm P}}V)(x,a)=r^{a}(x)+\gamma{\rm P}^{a}V(x)$ and the maximum selection operator $\mathsf{M}:\mathcal{B}(\mathsf{X}\times\mathsf{A})\rightarrow\mathcal{B}(\mathsf{X})$ is defined by $(\mathsf{M}V^{\cdot})(x)=\max_{a}V^{a}(x)$. Then $\mathsf{M}\mathsf{T}_{{\rm P}}$ corresponds to the Bellman optimality operator, see Puterman (2014). The optimal value function $V^{\star}$ satisfies a non-linear fixed-point equation
${V^{\star}(x)=\mathsf{M}\mathsf{T}_{{\rm P}}V^{\star}(x).}$ (1)
@@
which is known as the Bellman optimality equation. We write $Y^{x,a},x\in\mathsf{X},a\in\mathsf{A}$ for a random variable generated according to ${\rm P}^{a}(\cdot|x)$, and define a random Bellman operator $(\widetilde{\mathsf{T}}_{{\rm P}}V)(x)\mapsto r^{a}(x)+\gamma V(Y^{x,a})$. We say that a (deterministic) policy $\pi$ is greedy w.r.t. a function $V\in\mathcal{B}(\mathsf{X})$ if, for all $x\in\mathsf{X}$
@@ 3 Upper solutions and the main concept of UVIP
A straightforward approach to bound the policy error $\Delta_{\pi}(x)$ requires the estimation of the optimal value function $V^{\star}(x)$. Recall that $V^{\star}$ is a solution of the Bellman optimality equation (1). If the transition kernel $({\rm P}_{a})_{a\in\mathsf{A}}$ is known, the standard solution is the value iteration algorithm, see Bertsekas & Shreve (1978). In this algorithm, the estimates are recursively constructed via $V_{k+1}=\mathsf{M}\mathsf{T}_{{\rm P}}V_{k}$. Due to the Banach’s fixed point theorem, $\|V_{k}-V^{\star}\|_{\mathsf{X}}\leq\gamma^{k}\|V_{0}-V^{\star}\|_{\mathsf{X}},$ provided that $V_{0}\in\mathcal{B}(\mathsf{X})$. Moreover, $V_{k}(x)\geq V^{\star}(x)$ for any $x\in\mathsf{X}$ and $k\in\mathbb{N}$, provided that $V_{0}(x)\geq V^{\star}(x)$. For example, if $\|r^{a}\|_{\mathsf{X}}\leq R_{\max}$ for all $a\in\mathsf{A}$, we can take $V_{0}(x)=R_{\max}/(1-\gamma)$.
Unfortunately, (1) does not allow to represent $V^{\star}$ as an expectation and to reduce the problem of estimating $V^{\star}$ to a stochastic approximation problem. Moreover, if $({\rm P}^{a})_{a\in\mathsf{A}}$ is replaced by its empirical estimate $\hat{{\rm P}}^{a}$ the desired upper biasness property $V_{k}(x)\geq V^{\star}(x)$ is lost. Some recent works (e.g. Efroni et al. (2019)) suggested a modification of the optimism-based approach applicable in case of unknown $({\rm P}^{a})_{a\in\mathsf{A}}$. Yet this modification contains an additional optimization step, which is unfeasible beyond the tabular state- and action space problems. Therefore the problem of constructing upper bounds for the optimal value function $V^{\star}$ and policy error remains open and highly relevant. Below we describe our approach, which is based on the following key assumptions:
@@
The key concept of our algorithm is an upper solution, introduced below.
Definition 3.1.
@@
Upper solutions can be used to build tight upper bounds for the optimal value function $V^{\star}.$ Let $\Phi\in\mathcal{B}(\mathsf{X})$ be a martingale function w.r.t. the operator ${\rm P}^{a}$, that is, ${\rm P}^{a}\Phi(x)=0$ for all $a\in\mathsf{A},x\in\mathsf{X}$. Define $V^{\rm{up}}$ as a solution to the following fixed point equation:
@@
Note that unlike the optimal state value function $V^{\star}$, the upper solution $V^{\rm{up}}$ is represented as an expectation, which allows us to use various stochastic approximation methods to compute $V^{\rm{up}}$. The Banach’s fixed-point theorem implies that for iterates
@@
$V_{k+1}^{\rm{up}}(x)=\mathsf{E}[\mathsf{M}\widetilde{\mathsf{T}}_{\rm P}(V_{k}^{\rm{up}}-\Phi^{x,\cdot}_{\pi})(x)]=\mathsf{E}\bigl{[}\max_{a}\{r^{a}(x)+\gamma(V_{k}^{\rm{up}}(Y^{x,a})-\Phi_{\pi}^{x,a}(Y^{x,a}))\}\bigr{]}$ (3)
with $V_{0}^{\rm{up}}\in\mathcal{B}(\mathsf{X})$. Algorithm 1 contains the pseudocode of the UVIP for MDPs with finite state and action spaces. Several generalizations are discussed in the next section.
@@
Further note that by taking $\Phi^{x,a}(y)\doteq V^{\star}(y)-({\rm P}^{a}V^{\star})(x)$, we get with probability $1:$
$V^{\star}(x)=(\mathsf{M}\widetilde{\mathsf{T}}_{{\rm P}}(V^{\star}-\Phi^{x,\cdot}))(x)=\max_{a}\{r^{a}(x)+\gamma(V^{\star}(Y^{x,a})-\Phi^{x,a}(Y^{x,a}))\},$ (4)
@@
The upper solutions can be used to evaluate the quality of the policies and to construct confidence intervals for $V^{\star}$. It is clear that
@@
for any $k\in\mathbb{N}$ and $x\in\mathsf{X}$, thus a policy $\pi$ can be evaluated by computing the difference $\Delta^{{\rm up}}_{\pi,k}(x)\doteq V_{k}^{\rm{up}}(x)-V^{\pi}(x)\geq\Delta_{\pi}(x)$. Representations (3) and (4) imply
$\left\|V_{k+1}^{{\rm{up}}}-V^{\star}\right\|_{\mathsf{X}}\leq\gamma\left\|V_{k}^{{\rm{up}}}-V^{\star}\right\|_{\mathsf{X}}+2\gamma\left\|V^{{\rm\pi}}-V^{\star}\right\|_{\mathsf{X}},\quad k\in\mathbb{N}\,.$
@@
As a result $\Delta_{\pi}^{{\rm up}}=0$ if $\pi=\pi^{\star}$ and the corresponding confidence intervals collapses into one point. Moreover, for a policy $\pi$ which is greedy w.r.t. an action-value function $Q^{\pi}(x,a)$, it holds that $V^{\pi}(x)\geq V^{\star}(x)-2(1-\gamma)^{-1}\|Q^{\pi}-Q^{\star}\|_{\mathsf{X}\times\mathsf{A}}$ (see Szepesvári (2010)). Thus we can rewrite the bound (5) in terms of action-value functions
@@
The quantity $\Delta_{\pi,k}^{{\rm up}}$ can be used to measure the quality of policies $\pi$ obtained by many well-known algorithms like Reinforce (Williams (1992)), API (Bertsekas & Tsitsiklis (1996)), A2C (Mnih et al. (2016)) and DQN (Mnih et al. (2013), Mnih et al. (2015)).

### v5

@@ Abstract
Policy evaluation is an important instrument for the comparison of different algorithms in Reinforcement Learning (RL). However, even a precise knowledge of the value function $V^{\pi}$ corresponding to a policy $\pi$ does not provide reliable information on how far the policy $\pi$ is from the optimal one. We present a novel model-free upper value iteration procedure (UVIP) that allows us to estimate the suboptimality gap $V^{\star}(x)-V^{\pi}(x)$ from above and to construct confidence intervals for $V^{\star}$. Our approach relies on upper bounds to the solution of the Bellman optimality equation via the martingale approach. We provide theoretical guarantees for UVIP under general assumptions and illustrate its performance on a number of benchmark RL problems.
Communicated by Alexander Vladimirovich Gasnikov.
journal: JOTA
@@ 1 Introduction
The key objective of Reinforcement Learning (RL) is to learn an optimal agent’s behavior in an unknown environment. A natural performance metric is given by the value function $V^{\pi}$, which is the expected total reward of the agent following $\pi$. There are efficient algorithms to evaluate this quantity, e.g., temporal difference methods Sutton (1988), Tsitsiklis and Van Roy (1996). Unfortunately, even a precise knowledge of $V^{\pi}$ does not provide reliable information on how far the policy $\pi$ is from the optimal one. At the same time, practitioners are often interested in quantitative guarantees on the suboptimality gap (policy error) $\Delta_{\pi}(x)\doteq V^{\star}(x)-V^{\pi}(x)$ or, more generally, in tight confidence bounds for the optimal value function $V^{\star}$. To address this issue, a popular quality measure is the regret of the algorithm, which is the difference between the total sum of rewards accumulated when following the optimal policy and the sum of rewards obtained when following the current policy $\pi$ (see, e.g., Jaksch et al. (2010)). In the setting of finite state- and action-space Markov Decision Processes (MDP), there is a variety of regret bounds for popular RL algorithms like Q-learning Jin et al. (2018), optimistic value iteration Azar et al. (2017), and many others. Unfortunately, regret bounds beyond the discrete setup are much less common in the literature. An even more crucial drawback of the regret-based comparison is that regret bounds are typically pessimistic and rely on the unknown quantities of the underlying MDPs.
In this paper, we address the problem of estimating $\Delta_{\pi}(x)$ by constructing model-free upper confidence bounds for the optimal value function $V^{\star}$ and, consequently, for the policy error $V^{\star}-V^{\pi}$. Our starting point is the Bellman optimality equation, which characterizes $V^{\star}$ as a fixed point of the Bellman operator. Rather than trying to approximate $V^{\star}$ directly, we introduce the notion of an upper solution to the Bellman equation. By combining this idea with the martingale-based duality argument, we design an upper value iteration procedure (UVIP) which, given an arbitrary policy $\pi$, produces an almost-sure upper bound on $V^{\pi}$ and thus on $V^{\star}$ using only samples from the (unknown) transition kernel.
@@
  • We propose a novel approach to construct model-free confidence bounds for the optimal value function $V^{\star}$ based on a notion of upper solutions.
@@
  • We study convergence properties of the approximate UVIP in the case of general state and action spaces. In particular, we show that the variance of the resulting upper bound is small if $\pi$ is close to $\pi^{*}$, leading to tight confidence bounds for $V^{\star}$.
The paper is organized as follows. First, in Section 2, we briefly recall the main concepts related to MDPs and introduce some notations. Then, in Section 3, we discuss the contributions related to our paper. In Sections 4 and 5, we introduce the framework of UVIP and discuss its basic properties. In Section 6, we perform a theoretical study of the approximate UVIP. The numerical results are collected in Section 7. Section 8 concludes the paper. Section A in the appendix is devoted to the proof of the main theoretical results.
Notations and definitions.
For $N\in\mathbb{N}$, we define $N]\doteq\{1,\ldots,N\}$. Let us denote the space of bounded measurable functions with domain $\mathsf{X}$ by $\mathcal{B}(\mathsf{X})$, equipped with the norm $\|f\|_{\mathsf{X}}=\sup_{x\in\mathsf{X}}|f(x)|$ for any $f\in\mathcal{B}(\mathsf{X})$. In what follows, whenever a norm is uniquely identifiable from its argument, we will drop the index of the norm denoting the underlying space. We denote by ${\rm P}^{a}$ an $\mathcal{B}(\mathsf{X})\to\mathcal{B}(\mathsf{X})$ operator defined by $({\rm P}^{a}V)(x)=\int V(x^{\prime}){\rm P}^{a}(dx^{\prime}|x)$. For an arbitrary metric space $(\mathcal{X},\rho_{\mathcal{X}})$ and function $f:\mathcal{X}\to\mathbb{R}$, we define $\mathrm{Lip}_{\rho_{\mathcal{X}}}(f)\doteq\sup_{x\neq y}|f(x)-f(y)|/\rho_{\mathcal{X}}(x,y)$.
@@ 2 Preliminary
A Markov Decision Process (MDP) is a tuple $(\mathsf{X},\mathsf{A},\mathcal{P},r)$, where $\mathsf{X}$ is the state space, $\mathsf{A}$ is the action space, $\mathcal{P}=({\rm P}^{a})_{a\in\mathsf{A}}$ is the transition probability kernel, and $r=(r^{a})_{a\in\mathsf{A}}$ is the reward function. For each state $x\in\mathsf{X}$ and action $a\in\mathsf{A}$, ${\rm P}^{a}(\cdot|x)$ stands for a distribution over the states in $\mathsf{X}$, that is, the distribution over the next states given that action $a$ is taken in the state $x$. For each action $a\in\mathsf{A}$ and state $x\in\mathsf{X}$, $r^{a}(x)$ gives a reward received when action $a$ is taken in state $x$. An MDP describes the interaction of an agent and its environment. When an action $A_{t}\in\mathsf{A}$ at time $t$ is chosen by the agent, the state $X_{t}$ transitions to $X_{t+1}\sim{\rm P}^{A_{t}}(\cdot|X_{t})$. The agent’s goal is to maximize the expected total discounted reward, $\mathsf{E}[\sum_{t=0}^{\infty}\gamma^{t}r^{A_{t}}(X_{t})]$, where $0<\gamma<1$ is the discount factor. A rule describing the way an agent acts given its past actions and observations is called a policy. The value function of a policy $\pi$ in a state $x\in\mathsf{X}$, denoted by $V^{\pi}(x)$, is $V^{\pi}(x)=\mathsf{E}[\sum_{t=0}^{\infty}\gamma^{t}r^{A_{t}}(X_{t})|X_{0}=x]$, that is, the expected total discounted reward when the initial state is $X_{0}=x$, assuming the agent follows the policy $\pi$. Similarly, we define the action-value function $Q^{\pi}(x,a)=\mathsf{E}[\sum_{t=0}^{\infty}\gamma^{t}r^{A_{t}}(X_{t})|X_{0}=x,A_{0}=a]$. An optimal policy is one that achieves the maximum possible value amongst all policies in each state $x\in\mathsf{X}$. The optimal value for state $x$ is denoted by $V^{\star}(x)$. A deterministic Markov policy can be identified with a map $\pi:\mathsf{X}\to\mathsf{A}$, and the space of measurable deterministic Markov policies will be denoted by $\Pi$. When, in addition, the reward function is bounded, which we assume from now on, all the value functions are bounded, and one can always find a deterministic Markov policy that is optimal Puterman (2014). We also define a greedy policy w.r.t. the action-value function $Q(x,a)$, which is a deterministic policy $\pi(x)\in\operatorname*{argmax}_{a\in\mathsf{A}}Q(x,a)$. The Bellman return operator w.r.t. $\rm P$}, $\mathsf{T}_{{\rm P}}:\mathcal{B}(\mathsf{X})\to\mathcal{B}(\mathsf{X}\times\mathsf{A})$, is defined by $(\mathsf{T}_{{\rm P}}V)(x,a)=r^{a}(x)+\gamma{\rm P}^{a}V(x)$, and the maximum selection operator $\mathsf{M}:\mathcal{B}(\mathsf{X}\times\mathsf{A})\rightarrow\mathcal{B}(\mathsf{X})$ is defined by $(\mathsf{M}V^{\cdot})(x)=\max_{a}V^{a}(x)$. Then $\mathsf{M}\mathsf{T}_{{\rm P}}$ corresponds to the Bellman optimality operator; see Puterman (2014). The optimal value function $V^{\star}$ satisfies a non-linear fixed-point equation
$V^{\star}(x)=\mathsf{M}\mathsf{T}_{{\rm P}}V^{\star}(x),$ (1)
@@
which is known as the Bellman optimality equation. We write $Y^{x,a}$, $x\in\mathsf{X}$, $a\in\mathsf{A}$, for a random variable generated according to ${\rm P}^{a}(\cdot|x)$ and define a random Bellman operator $(\widetilde{\mathsf{T}}_{{\rm P}}V)(x)\mapsto r^{a}(x)+\gamma V(Y^{x,a})$. We say that a (deterministic) policy $\pi$ is greedy w.r.t. a function $V\in\mathcal{B}(\mathsf{X})$ if, for all $x\in\mathsf{X}$,
@@
3 Related Works
There is a large body of work on theoretical guarantees for $\Delta_{\pi}(x)$ in approximate dynamic programming and model-based RL, including results on fitted Value- and $Q$-iteration and on policy error bounds for model-based approaches with factored linear models; see, for example, Antos et al. (2007), Szepesvári (2010), Ávila Pires and Szepesvári (2016) and references therein. These bounds typically depend on the problem characteristics, which are not known in practice. Moreover, they are often tied to the specific algorithm that produced $\pi$ and are not directly applicable as a generic evaluation tool for an arbitrary policy. For instance, in Approximate Policy Iteration (API, Bertsekas and Tsitsiklis (1996)), all existing bounds for $\Delta_{\pi}(x)$ depend on the one-step error induced by the approximation of the $Q$-function. This one-step error is difficult to quantify since it depends on the unknown smoothness properties of the $Q$-function. Similarly, in policy gradient methods (see, e.g., Sutton and Barto (2018)), there is an approximation error due to the choice of the family of policies that can be hardly quantified.
The approach based on the policy optimism principle (see Efroni et al. (2019)) suggests to initialise the value iteration algorithm using an upper bound (optimistic value) for $V^{\star}$, yielding a sequence of upper bounds converging to $V^{\star}$. Similarly, UCRL2 and its refinements provide near-minimax regret guarantees in the average-reward and finite-horizon settings; see Jaksch et al. (2010), Azar et al. (2017) and Bourel et al. (2020). However, these approaches are tailored to finite state- and action-space MDPs and are not applicable to evaluate the quality of a general policy $\pi$.
The concept of upper solutions is closely related to martingale duality in optimal control and the information relaxation approach; see Belomestny and Schoenmakers (2018), Rogers (2007) and Brown and Smith (2022). This idea has been successfully used in the recent paper Shar and Jiang (2020). This work proposes to use the duality approach to improve the performance of the $Q$-learning algorithm in finite-horizon MDPs through the use of “lookahead” upper and lower bounds. In the subsequent work El Shar and Jiang (2023), the authors extend their duality-based ideas to structured weakly coupled MDPs. Furthermore, in Chen et al. (2025), the authors propose a duality-based algorithm, ADRL, which utilizes neural network approximation for high-dimensional stochastic control problems. There are also recent papers focusing on lookahead-based methods Rosenberg et al. (2023), Merlis (2024). Specifically, in Rosenberg et al. (2023), the authors propose adaptive planning horizons for planning and deep $Q$-learning, choosing the depth of lookahead as a function of state-dependent value estimates. In Merlis (2024), the authors propose a regret-optimal algorithm where the agent receives additional stochastic lookahead information (e.g., transition or reward realizations before acting).
The concept of upper solutions also has a connection to distributional RL, as it can be formulated pathwise or using the distributional Bellman operator; see, e.g., Lyle et al. (2019). Development of the distributional counterpart of the upper solution to the Bellman equation is a promising future research area.
@@ 4 Upper Solutions and the Main Concept of UVIP
A straightforward approach to bound the policy error $\Delta_{\pi}(x)$ requires the estimation of the optimal value function $V^{\star}(x)$. Recall that $V^{\star}$ is a solution of the Bellman optimality equation (1). If the transition kernel $({\rm P}_{a})_{a\in\mathsf{A}}$ is known, the standard solution is the value iteration algorithm; see Bertsekas and Shreve (1996). In this algorithm, the estimates are constructed recursively via $V_{k+1}=\mathsf{M}\mathsf{T}_{{\rm P}}V_{k}$. Due to Banach’s fixed point theorem, $\|V_{k}-V^{\star}\|_{\mathsf{X}}\leq\gamma^{k}\|V_{0}-V^{\star}\|_{\mathsf{X}},$ provided that $V_{0}\in\mathcal{B}(\mathsf{X})$. Moreover, $V_{k}(x)\geq V^{\star}(x)$ for any $x\in\mathsf{X}$ and $k\in\mathbb{N}$, provided that $V_{0}(x)\geq V^{\star}(x)$. For example, if $\|r^{a}\|_{\mathsf{X}}\leq R_{\max}$ for all $a\in\mathsf{A}$, we can take $V_{0}(x)=R_{\max}/(1-\gamma)$.
Unfortunately, (1) does not allow us to represent $V^{\star}$ as an expectation and to reduce the problem of estimating $V^{\star}$ to a stochastic approximation problem. Moreover, if $({\rm P}^{a})_{a\in\mathsf{A}}$ is replaced by its empirical estimate $\hat{{\rm P}}^{a}$, the desired upper-bias property $V_{k}(x)\geq V^{\star}(x)$ is lost. Some recent work (e.g., Efroni et al. (2019)) suggested a modification of the optimism-based approach applicable in case of unknown $({\rm P}^{a})_{a\in\mathsf{A}}$. Yet this modification contains an additional optimization step, which is unfeasible beyond tabular state- and action-space problems. Therefore, the problem of constructing upper bounds for the optimal value function $V^{\star}$ and the policy error remains open and highly relevant. In the following, we describe our approach, which is based on the following key assumptions:
@@
The key concept of our algorithm is upper solution, introduced below.
Definition 4.1
@@
Upper solutions can be used to build tight upper bounds for the optimal value function $V^{\star}.$ Let $\Phi^{x,a}\in\mathcal{B}(\mathsf{X})$, $x\in\mathsf{X}$, $a\in\mathsf{A}$, be a family of martingale functions w.r.t. the operator ${\rm P}^{a}$, that is, ${\rm P}^{a}\Phi^{x,a}(x)=0$ for all $a\in\mathsf{A},x\in\mathsf{X}$. Define $V^{\rm{up}}$ as a solution to the following fixed-point equation:
@@
Note that unlike the optimal state value function $V^{\star}$, the upper solution $V^{\rm{up}}$ is represented as an expectation, which allows us to use various stochastic approximation methods to compute $V^{\rm{up}}$. Banach’s fixed-point theorem implies that for iterates
@@
[…]
with $V_{0}^{\rm{up}}\in\mathcal{B}(\mathsf{X})$. The algorithm 1 contains the pseudocode of the UVIP for MDPs with finite state and action spaces. Several generalizations are discussed in the next section.
@@
Taking $\Phi^{x,a}(y)\doteq V^{\star}(y)-({\rm P}^{a}V^{\star})(x)$, we get with probability $1$:
[…]
@@
The upper solutions can be used to evaluate the quality of policies and to construct confidence intervals for $V^{\star}$. It is clear that
@@
for any $k\in\mathbb{N}$ and $x\in\mathsf{X}$, and thus a policy $\pi$ can be evaluated by computing the difference $\Delta^{{\rm up}}_{\pi,k}(x)\doteq V_{k}^{\rm{up}}(x)-V^{\pi}(x)\geq\Delta_{\pi}(x)$. Representations (3) and (4) imply
$\left\|V_{k+1}^{{\rm{up}}}-V^{\star}\right\|_{\mathsf{X}}\leq\gamma\left\|V_{k}^{{\rm{up}}}-V^{\star}\right\|_{\mathsf{X}}+2\gamma\left\|V^{\pi}-V^{\star}\right\|_{\mathsf{X}},\quad k\in\mathbb{N}\,.$
@@
As a result, $\Delta_{\pi}^{{\rm up}}=0$ if $\pi=\pi^{\star}$ and the corresponding confidence intervals collapse into a single point. Moreover, for a policy $\pi$ which is greedy w.r.t. an action-value function $Q^{\pi}(x,a)$, it holds that $V^{\pi}(x)\geq V^{\star}(x)-2(1-\gamma)^{-1}\|Q^{\pi}-Q^{\star}\|_{\mathsf{X}\times\mathsf{A}}$ (see Szepesvári (2010)). Thus, we can rewrite the bound (5) in terms of action-value functions
@@
The quantity $\Delta_{\pi,k}^{{\rm up}}$ can be used to measure the quality of policies $\pi$ obtained by many well-known algorithms like Reinforce (Williams (1992)), API (Bertsekas and Tsitsiklis (1996)), A2C (Mnih et al. (2016)) and DQN (Mnih et al. (2015)).
Comparison with PAC-based confidence intervals and IPOC framework.
We highlight the core differences between our approach and PAC-based confidence intervals. Typically, the latter provide instance-specific bounds (depending on the particular policy $\pi$ and on the properties of the particular algorithm that outputs $\pi$), which additionally depend on problem characteristics that are practically unknown. In contrast, we aim to suggest a generic approach to estimating $\Delta^{\pi}(x)$ for an arbitrary input policy $\pi$. Our approach can be integrated into the IPOC framework Dann et al. (2019), since it provides a suboptimality gap that can be interpreted precisely as an optimality certificate. At the same time, the IPOC approach is not a direct counterpart of the $\sf{UVIP}$ procedure, as IPOC itself does not provide an estimate of the suboptimality gap $\Delta^{\pi}(x)$ (the optimality certificate in the terminology of Dann et al. (2019)), but instead relies on estimates derived from PAC-style analyses of particular policies $\pi$. Moreover, following the analysis in Dann et al. (2019), one can translate the finite-sample bounds on the $\sf{UVIP}$ error given in Theorem 6.1 into regret bounds. We leave the detailed analysis of the IPOC procedure based on $\sf{UVIP}$ outputs for particular MDP settings as an important direction for future work.